How many grams of Fe3O4 are needed to react with 8.8 x1022 molecules of CO?
1
Expert's answer
2013-02-25T11:14:12-0500
The chemical equation for this reaction isFe3O4 + 4CO --> 3Fe + 4CO2

According to Avogadro law: n(moles) = N/Na ( N - the number of
molecules, Na - 6.02 x 10^23)

n(CO) = (8.8 x 10^22)/(6.02 x 10^23) = 0.146 moles

According to equation n(Fe3O4) = n(CO)/4

n(Fe3O4) = 0.146/4 = 0.0365 moles

m(grams) = MW x n(moles)

MW(Fe3O4) = 3 x MW(Fe) + 4 x MW(O) = 3 x 55.8 + 4 x 16.0 = 231.4 grams/mol

m(Fe3O4) = 231.4 x 0.0365 = 8.45 grams

Answer: m(Fe3O4) = 8.45 grams